40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9252     Entry time: Thu Oct 17 22:38:25 2013
Author: Masayuki 
Type: Update 
Category: PSL 
Subject: PSL temperature changed 

[Manasa, Masayuki]

PSL temperature changed

The beat note of Xarm looked somehow strange before (elog). It should be the highest when the green transmission power is highest, in other words when the end green PDH locks with a TEM00 mode. But it was not like that. When the end PDH locked with other modes (GTRX: below 0.3), the beat note was higher than TEM00 mode (GTRX: around 0.5). 

We guessed that end green laser was somewhere around the point where there were 2 stable TEM00 modes . In order to move away from this unstable region of the end laser, we changed PSL temperature to obtain beat note at a different green laser frequency where we do not have any of the weird modes oscillating.

We changed the PSL temperature from 31.63 degree to 31.33 degree. We measured the in-loop noise of ALS loop and I attached it. There is not big difference in Yarm, but the Xarm in-loop noise become better in high frequency region. We think before the xend green laser was in a not-so-good state and the laser had more frequency noise then. 

ALS stability

After change PSL temperature, Xarm ALS is so stable. Actually Xarm is being locked right now and it is locked for more than 50 minutes!!
Although the Xarm ALS is so stable, Yarm ALS is not stable right now. It lost lock by ~5min. We don't know what is the reason, so we will try to fix it tomorrow.

Attachment 1: in-loop.pdf  72 kB  Uploaded Thu Oct 17 23:39:24 2013  | Hide | Hide all
in-loop.pdf
ELOG V3.1.3-